# Non-Expansive Mappings in Two-Time-Scale Stochastic Approximation: Finite-Time Analysis

### v1

@@ Abstract
Two-time-scale stochastic approximation is an iterative algorithm used in applications such as optimization, reinforcement learning, and control. Finite-time analysis of these algorithms has primarily focused on fixed point iterations where both time-scales have contractive mappings. In this paper, we study two-time-scale iterations, where the slower time-scale has a non-expansive mapping. For such algorithms, the slower time-scale can be considered a stochastic inexact Krasnoselskii-Mann iteration. We show that the mean square error decays at a rate $\mathcal{O}(1/k^{1/4-\epsilon})$, where $\epsilon>0$ is arbitrarily small. We also show almost sure convergence of iterates to the set of fixed points. We show the applicability of our framework by applying our results to minimax optimization, linear stochastic approximation, and Lagrangian optimization.
@@ 1 Introduction
The majority of analysis for two-time-scale algorithms has occurred under the assumption that the function $f(x,y)$ is contractive in $x$, and the function $g(x^{*}(y),y)$ is contractive [10, 18, 2]. Under these assumptions, the algorithm converges to $x^{*},y^{*}$ where they are the unique solution to the equations: $f(x,y)=x$ and $g(x,y)=y$ [2]. In recent years, there have been works that obtain finite-time bounds for iterates, but they also mainly focus on contractive mappings in both time-scales [10, 18, 16].
@@
    We present a finite-time mean square bound for our algorithm, which has contractive and non-expansive mappings in the faster and slower time-scales, respectively. We show that the mean square error decays at a rate $\mathcal{O}(k^{-1/4+\epsilon})$, where $\epsilon>0$ can be arbitrarily small depending on the step size.
@@
    Using similar techniques, we present an additional result where the slower time-scale iteration contains a gradient descent operator for a smooth function. The gradient descent operator is not necessarily non-expansive in this case. In this setting, we show that the minimum, from time 00 to $k$, of the expected norm of the gradient is bounded by $\mathcal{O}(\log(k)/k^{2/5})$.
@@ 1.1 Related Work
Two-time-scale algorithms with contractive mappings (in both time-scales) play an important role in RL [27], and hence finite time bounds for such algorithms have also been studied in detail. In this setting, the bound is obtained on $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}]$. A major focus in this field is on linear two-time-scale SA, i.e., where the maps $f(\cdot)$ and $g(\cdot)$ are linear. In [18], an asymptotic convergence rate of $\mathcal{O}(1/k)$ is obtained, while in [16, 17] the result is extended to a finite-time bound. We also study linear two-time-scale SA as a special case. But we study the case where the matrix relevant for the slower time-scale is negative semidefinite instead of Hurwitz or negative definite as assumed in the above works.
Bounds for contractive non-linear two-time-scale SA have also been studied in recent years. In [10], a rate of $\mathcal{O}(1/k^{2/3})$ is obtained for this setting. In [11], a rate of $\mathcal{O}(1/k)$ is obtained but a time-weighted average of iterates is considered.
@@ 1.2 Outline and Notation
The paper is structured as follows: Section 2 sets up and presents the main result. The projected variant is presented in Subsection 2.1. Section 3 contains a proof sketch for our main result. Section 4 presents the framework where the slower time-scale has a gradient descent operator. Section 5 presents the above-listed applications and how they can be incorporated into our framework. Finally, Section 6 concludes the paper and presents some future directions. Appendices A-D contain proofs for our results.
[…]
@@ 2 Non-Expansive Mappings in Slower Time-Scale
We first make the assumption that function $f(x,y)$ is contractive in $x$. {assumption} The function $f:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{1}}$ is contractive in $x$ for any given $y\in\mathbb{R}^{d_{2}}$, i.e.,
@@
for all $x_{1},x_{2}\in\mathbb{R}^{d_{1}}$ and $y\in\mathbb{R}^{d_{2}}$. Here $\mu<1$ is the contraction factor. By the Banach contraction mapping theorem, the above assumption implies that $f(\cdot,y)$ has a unique fixed point for each $y$. That is, for each $y\in\mathbb{R}^{d_{2}}$, there exists unique $x^{*}(y)$ such that $f(x^{*}(y),y)=x^{*}(y)$. The following is an assumption on the function $g:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{2}}$. {assumption} The function $g(x^{*}(\cdot),\cdot):\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{2}}$ is non-expansive, i.e.,
@@
The opposite directions are also true as shown in Lemma D.1 in Appendix D.
@@
$\alpha_{k}=\frac{\alpha_{0}}{(k+1)^{\mathfrak{a}}}\;\;\text{and}\;\;\beta_{k}=\frac{\beta_{0}}{(k+1)^{\mathfrak{b}}},$
where $10.5<\mathfrak{a}<\mathfrak{b}<$ 1 and $\alpha_{0},\beta_{0}\leq 0.5$. This ensures that they are non-summable $(\sum_{k}\alpha_{k}=\infty,\sum_{k}\beta_{k}=\infty)$, square-summable $(\sum_{k}\alpha^{2}_{k}<\infty,\sum_{k}\beta^{2}_{k}<\infty)$ and non-increasing $(\alpha_{k+1}\leq\alpha_{k}$ and $\beta_{k+1}\leq\beta_{k})$. This also ensures that $\lim_{k\uparrow\infty}\beta_{k}/\alpha_{k}=0$. Finally, we make the key assumption that $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1.$ We restrict our attention to step sizes of this form for simplicity but our results can be extended to other step size sequences. We do not include $1/(k+1)$ in our set of step sizes because it is not optimal for non-expansive fixed point iterations. The assumptions that the step sizes are non-summable, square summable, and satisfy $\beta_{k}/\alpha_{k}$ goes to zero as $k\uparrow\infty$ are standard in stochastic approximation literature [2]. The assumption of $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1$ is the assumption that decides the time-scale separation between the updates and is key for our analysis.
We are now ready to state our main result. An outline of the proof is given in Section 3, and the complete proof is in Appendix A.
@@
Suppose that Assumptions 2-2 hold. Let $\{x_{k},y_{k}\}$ be generated by (1) with $\beta_{0}/\alpha_{0}\leq(1-\mu)/(2L^{2})$. Then,
@@
Optimal Rate and Step size Choice
The optimal rate for the algorithm based on the above analysis is $1/(k+1)^{0.25-\epsilon}$ where $\epsilon>0$ can be arbitrarily small. This can be achieved by choosing $\mathfrak{a}=0.5+(2/3)\epsilon$ and $\mathfrak{b}=0.75+\epsilon$. Unlike contractive maps, in the analysis of fixed point iterations with non-expansive mappings, a slowly decreasing step size gives a better rate. Hence, a slowly decreasing $\beta_{k}$, or equivalently a smaller $\mathfrak{b}$, gives a better rate. This, along with the assumptions that $\mathfrak{a}>0.5$ and $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1$ (equivalently, $2\mathfrak{b}\geq 3\mathfrak{a}$), dictates the optimal rate.
The assumptions on $\beta_{0}/\alpha_{0}$ are required only for the bounds to hold for all $k\geq 0$. Instead, if we have that $\beta_{k}/\alpha_{k}\leq(1-\mu)/2L^{2}$ holds for all $k\geq K$, then the bounds will hold for all $k\geq K$. The existence of such a $K$ is guaranteed by Assumption 2.
We also want to note that $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$ and $\mathbb{E}\left[\|g(x^{*}(y_{k}),y_{k})-y_{k}\|^{2}\right]$ are $\mathcal{O}(\alpha_{k})$ and $\mathcal{O}(1/(k\beta_{k}))$, respectively.
@@ 2.1 Projected Variant
We still assume that the function $f(\cdot,y)$ is contractive for all $y$. But the relevant fixed point is now the fixed point for $\mathcal{P}_{\mathcal{X}}(f(\cdot,y))$, i.e., we now define $\hat{x}(y)$ as the unique point that satisfies $\hat{x}(y)=\mathcal{P}_{\mathcal{X}}(f(\hat{x}(y),y))$. In this case, the function $g(\hat{x}(\cdot),\cdot)$ is assumed to be non-expansive. {assumption} The function $g(\hat{x}(\cdot),\cdot):\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{2}}$ is non-expansive, i.e.,
@@
for all $y_{1},y_{2}\in\mathbb{R}^{d_{2}}$. Let us denote the set of fixed points of $g(\hat{x}(\cdot),\cdot)$ by $\mathcal{\hat{Y}}$, i.e., $\mathcal{\hat{Y}}=\{y\in\mathbb{R}^{d_{2}}\mid g(\hat{x}(y),y)=y\}$. We assume that the set $\mathcal{\hat{Y}}$ is non-empty.
This projected variant is studied as a special case because there exist scenarios where the projection of the faster time-scale updates is why the slower time-scale iteration has a non-expansive mapping (instead of a contractive mapping). Formally, there arise scenarios where the map $g(x^{*}(\cdot),\cdot)$ is contractive, but the map $g(\hat{x}(\cdot),\cdot)$ is not guaranteed to be contractive, and is only non-expansive. We study one such scenario in Subsection 5.3.
The main theorem for this projected variant is similar to Theorem 2.1, and its proof is presented in Appendix B.
Theorem 2.2.
Suppose that Assumptions 2, 2.1 and 2-2 hold. Let $\{x_{k},y_{k}\}$ be generated by (2) with $\beta_{0}/\alpha_{0}\leq(1-\mu)/(2L^{2})$. Then,
@@ 3 Proof Outline
The assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}$ is key for obtaining the above intermediate bounds. We first add the two intermediate bounds. Then we show that the iterates are bounded in expectation using the assumption that the step size sequence $\alpha_{k}$ is square-summable.
@@
    For some $\Gamma_{7}>0$,

### v3

@@ Abstract
Two-time-scale stochastic approximation algorithms are iterative methods used in applications such as optimization, reinforcement learning, and control. Finite-time analysis of these algorithms has primarily focused on fixed point iterations where both time-scales have contractive mappings. In this work, we broaden the scope of such analyses by considering settings where the slower time-scale has a non-expansive mapping. For such algorithms, the slower time-scale can be viewed as a stochastic inexact Krasnoselskii-Mann iteration. We also study a variant where the faster time-scale has a projection step which leads to non-expansiveness in the slower time-scale. We show that the last-iterate mean square residual error for such algorithms decays at a rate $\mathcal{O}(1/k^{1/4-\epsilon})$, where $\epsilon>0$ is arbitrarily small. We further establish almost sure convergence of iterates to the set of fixed points. We demonstrate the applicability of our framework by applying our results to minimax optimization, linear stochastic approximation, and Lagrangian optimization.
@@ 1 Introduction
The majority of analysis for two-time-scale algorithms has occurred under the assumption that the function $f(x,y)$ is contractive in $x$, and the function $g(x^{*}(y),y)$ is contractive [5, 16, 22]. Under these assumptions, the algorithm converges to $x^{*},y^{*}$ where they are the unique solution to the equations: $f(x,y)=x$ and $g(x,y)=y$ [2]. In recent years, there have been works that obtain finite-time bounds for iterates, but they also mainly focus on contractive mappings in both time-scales [5, 16].
@@
    We present a finite-time bound for our algorithm, which has contractive and non-expansive mappings in the faster and slower time-scales, respectively. We show that the last-iterate mean square residual error decays at a rate $\mathcal{O}(k^{-1/4+\epsilon})$, where $\epsilon>0$ can be arbitrarily small depending on the step size.
@@
    Using similar techniques, we present an additional result where the slower time-scale iteration contains a gradient descent operator for a smooth function. The gradient descent operator is not necessarily non-expansive in this case. In this setting, we show that the minimum, from time 0 to $k$, of the expected norm of the gradient is bounded by $\mathcal{O}(\log(k)/k^{2/5})$.
@@ 1.1 Related Work
Two-time-scale algorithms with contractive mappings (in both time-scales) play an important role in RL [32], and hence finite time bounds for such algorithms have also been studied in detail. In this setting, the bound is obtained on $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}]$. A major focus in this field is on linear two-time-scale SA, i.e., where the maps $f(\cdot)$ and $g(\cdot)$ are linear. In [22], an asymptotic convergence rate of $\mathcal{O}(1/k)$ is obtained, while in [13, 14, 20, 21] the result is extended to a finite-time bound. We also study linear two-time-scale SA as a special case. But we study the case where the matrix relevant for the slower time-scale is negative semidefinite instead of Hurwitz or negative definite as assumed in the above works.
Bounds for contractive non-linear two-time-scale SA have also been studied in recent years. A rate of $\mathcal{O}(1/k^{2/3})$ was previously established for this setting in [16]. This rate has since been improved to $\mathcal{O}(1/k)$ in [5]. Rates for two-time-scale SA with contractions under arbitrary norms and Markovian noise were obtained in [8].
@@ 1.2 Outline and Notation
The paper is structured as follows: Section 2 sets up and presents the main result. The projected variant is presented in Subsection 2.1. Section 3 contains a proof sketch for our main result. Section 4 presents the framework where the slower time-scale has a gradient descent operator. Section 5 presents the above-listed applications and how they can be incorporated into our framework. Finally, Section 7 concludes the paper and presents some future directions. Appendices A-B contain proofs for our results.
Throughout this work, $\|\cdot\|$ denotes the Euclidean norm, and $\langle x_{1},x_{2}\rangle$ denotes the inner product given by $x_{1}^{T}x_{2}$.
@@ 2 Non-Expansive Mappings in Slower Time-Scale
We first make the assumption that function $f(x,y)$ is contractive in $x$. {assumption} The function $f:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{1}}$ is contractive in $x$ for each $y\in\mathbb{R}^{d_{2}}$. Specifically, there exists a constant $0\leq\mu<1$, independent of $y$, such that
@@
for all $x_{1},x_{2}\in\mathbb{R}^{d_{1}}$ and $y\in\mathbb{R}^{d_{2}}$. By the Banach contraction mapping theorem, the above assumption implies that $f(\cdot,y)$ has a unique fixed point for each $y$. That is, for each $y\in\mathbb{R}^{d_{2}}$, there exists unique $x^{*}(y)$ such that $f(x^{*}(y),y)=x^{*}(y)$. The following is an assumption on the function $g:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{2}}$. {assumption} The function $g(x^{*}(\cdot),\cdot):\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{2}}$ is non-expansive, i.e.,
@@
For completeness, the complete equivalence, including the proofs for the above statements and the converse directions, is presented in Appendix E.
@@
$\alpha_{k}=\frac{\alpha}{(k+K_{1})^{\mathfrak{a}}}\;\;\text{and}\;\;\beta_{k}=\frac{\beta}{(k+K_{1})^{\mathfrak{b}}},$
where $0.5<\mathfrak{a}<\mathfrak{b}<1$, $\alpha,\beta\leq 0.5$ and $K_{1}\geq 1$. This ensures that they are non-summable $(\sum_{k}\alpha_{k}=\infty,\sum_{k}\beta_{k}=\infty)$, square-summable $(\sum_{k}\alpha^{2}_{k}<\infty,\sum_{k}\beta^{2}_{k}<\infty)$ and non-increasing $(\alpha_{k+1}\leq\alpha_{k}$ and $\beta_{k+1}\leq\beta_{k})$. This also ensures that $\lim_{k\uparrow\infty}\beta_{k}/\alpha_{k}=0$. Finally, we make the key assumption that $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1.$ We restrict our attention to step sizes of this form for simplicity but our results can be extended to other step size sequences. We do not include $1/(k+K_{1})$ in our set of step sizes because it is not optimal for non-expansive fixed point iterations. The assumptions that the step sizes are non-summable, square summable, and satisfy $\beta_{k}/\alpha_{k}$ goes to zero as $k\uparrow\infty$ are standard in stochastic approximation literature [2]. The assumption of $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1$ is the assumption that decides the time-scale separation between the updates. The assumption is key for our analysis as it is the deciding factor in the bound on the mean square error. It is required to show that the iterates are bounded in expectation (Lemma 3.3) which is a key step in the proof. Similar assumptions have also been taken in papers such as [16].
Define $\gamma_{1}\coloneqq(1-\mu)/(2L^{2})$ and $\gamma_{2}\coloneqq\max\left\{\left(2\mathfrak{a}/(\mu\alpha)\right)^{1/1-\mathfrak{a}},\left(2\mathfrak{b}/\beta\right)^{1/1-\mathfrak{b}}\right\}$. The following is our main result. An outline of its proof is given in Section 3, and the complete proof has been presented in Appendix A.
@@
Suppose that Assumptions 2-2 hold. Let $\{x_{k},y_{k}\}$ be generated by (1) with $\beta/\alpha\leq\gamma_{1}$ and $K_{1}\geq\gamma_{2}$. Then,
@@
Best Rate and Step size Choice
The best rate for the algorithm based on the above analysis is $1/(k+1)^{0.25-\epsilon}$ where $\epsilon>0$ can be arbitrarily small. This can be achieved by choosing $\mathfrak{a}=0.5+(2/3)\epsilon$ and $\mathfrak{b}=0.75+\epsilon$. Unlike contractive maps, in the analysis of fixed point iterations with non-expansive mappings, a slowly decreasing step size gives a better rate. Hence, a slowly decreasing $\beta_{k}$, or equivalently a smaller $\mathfrak{b}$, gives a better rate. This, along with the assumptions that $\mathfrak{a}>0.5$ and $\beta_{k}^{2}/\alpha_{k}^{3}\leq 1$ (equivalently, $2\mathfrak{b}\geq 3\mathfrak{a}$), dictates the optimal rate.
We also point out that $\mathbb{E}\!\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$ and $\mathbb{E}\!\left[\|g(x^{*}(y_{k}),y_{k})-y_{k}\|^{2}\right]$ scale as $\mathcal{O}(\alpha_{k})$ and $\mathcal{O}(1/(k\beta_{k}))$, respectively. For the given stepsize sequences $\{\alpha_{k}\}$ and $\{\beta_{k}\}$, these rates are the best known upper bounds for stochastic approximation with contractive and non-expansive operators, respectively, even in the single-time-scale setting [10, 3]. Consequently, when optimizing over all stepsize sequences admissible under Assumption 2, the overall rate is fundamentally constrained by the stepsize requirements, and could potentially be improved only if these constraints were relaxed.
The assumptions on $\beta/\alpha$ and $K_{1}$ are required only for the bounds to hold for all $k\geq 0$. Instead, if we have that $\beta_{k}/\alpha_{k}\leq\gamma_{1}$ for all $k\geq K_{0}$, then the bounds will hold for all $k\geq\max\{K_{0},\gamma_{2}\}$. The existence of such $K_{0}$ is guaranteed by Assumption 2.
Comparison with Rates for Contractive SA
The mean square error rate of $\mathcal{O}(k^{-1/4+\epsilon})$ is weaker than bounds observed for contractive SA. But such weaker bounds are frequently observed in non-expansive variants of SA. The following table compares bounds for contractive and non-expansive SA under different settings.
Table 1: Comparison of current best bounds for SA schemes under contractive and non-expansive mappings. The Euclidean and arbitrary norms refer to the norms under which the mapping satisfies contractive or non-expansive properties, respectively.
Setting Contractive Maps Non-Expansive Maps
Single Time-Scale SA
(Euclidean norm)
$\mathbb{E}\left[\|x_{k}-x^{*}\|^{2}\right]=\mathcal{O}\left(1/k\right)$
(Stepsize $\mathcal{O}(1/k)$) [10]
$\mathbb{E}\left[\|f(x_{k})-x_{k}\|^{2}\right]=\mathcal{O}(\log(k)/k^{1/3})$
(Stepsize $\mathcal{O}(1/k^{2/3})$) [3]
Single Time-Scale SA
(arbitrary norm)
$\mathbb{E}\left[\|x_{k}-x^{*}\|^{2}\right]=\mathcal{O}(1/k)$
(Stepsize $\mathcal{O}(1/k)$) [10]
$\mathbb{E}\left[\|f(x_{k})-x_{k}\|^{2}\right]=\mathcal{O}(\log(k)/k^{1/3})$
(Stepsize $\mathcal{O}(1/k^{2/3})$) [3]
Two-Time-Scale SA $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]=\mathcal{O}(1/k)$
$\mathbb{E}\left[\|y_{k}-y^{*}\|^{2}\right]=\mathcal{O}(1/k)$
(Stepsizes $\mathcal{O}(1/k)$) [5]
$\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]=\mathcal{O}(1/k^{0.5})$
$\mathbb{E}\left[\|g(x^{*}(y_{k}),y_{k})-y_{k}\|^{2}\right]=\mathcal{O}(1/k^{0.25-\epsilon})$
[Our result]
While a bound of $\mathcal{O}(1/k)$ is achieved in all the cases of contractive SA listed above, the bound achieved by their non-expansive analogs is much weaker. For a stepsize of the form $\mathcal{O}(1/k^{\mathfrak{b}})$, the rate achieved in non-expansive SA is usually of the order $\mathcal{O}(1/k^{1-\mathfrak{b}})$. As we require square-summability, we need $\mathfrak{b}>0.5$. The additional assumptions on the stepsize sequences further weaken the bound in our setting.
@@ 2.1 Projected Variant
We still assume that the function $f(\cdot,y)$ is contractive for all $y$. But the relevant fixed point is now the fixed point for $\mathcal{P}_{\mathcal{X}}(f(\cdot,y))$, i.e., we now define $\hat{x}(y)$ as the unique point that satisfies $\hat{x}(y)=\mathcal{P}_{\mathcal{X}}(f(\hat{x}(y),y))$. In this case, the function g(x^(),)g($\hat${x}($\cdot$$\cdot)$ is assumed to be non-expansive. {assumption} The function $g(\hat{x}(\cdot),\cdot):\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{2}}$ is non-expansive, i.e.,
@@
for all $y_{1},y_{2}\in\mathbb{R}^{d_{2}}$. Let us denote the set of fixed points of g(x^(),)g($\hat${x}($\cdot$$\cdot)$ by $\mathcal{\hat{Y}}$, i.e., $\mathcal{\hat{Y}}=\{y\in\mathbb{R}^{d_{2}}\mid g(\hat{x}(y),y)=y\}$. We assume that the set $\mathcal{\hat{Y}}$ is non-empty.
This projected variant is studied as a special case because there exist scenarios where the projection of the faster time-scale updates is why the slower time-scale iteration has a non-expansive mapping (instead of a contractive mapping). Formally, there arise scenarios where the map $g(x^{*}(\cdot),\cdot)$ is contractive, but the map g(x^(),)g($\hat${x}($\cdot$$\cdot)$ is not guaranteed to be contractive, and is only non-expansive. We study one such scenario in Subsection 5.3.
The main theorem for this projected variant is similar to Theorem 2.1, and its proof is presented in Appendix C.
Theorem $2.$
Suppose that Assumptions 2, 2.1 and 2-2 hold. Let $\{x_{k},y_{k}\}$ be generated by (2) with $\beta/\alpha\leq\gamma_{1}$ and $K_{1}\geq\gamma_{2}$. Then,
@@ 3 Proof Outline
When deriving the bound for the faster time-scale, we encounter an error term which has a dependence of $\beta_{k}^{2}/\alpha_{k}$ on $k$. The assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}$ helps us bound this term by $\alpha_{k}^{2}$. Since we have the assumption that $\alpha_{k}$ is square-summable, we get that this error term is summable. Hence the assumptions $\beta_{k}^{2}\leq\alpha_{k}^{3}$ and square-summability of $\alpha_{k}$ are key to show that the iterates are bounded in expectation.
@@
    For all $i\leq k-1$, there exists $\Gamma_{7}>0$ such that